OpenStudy (nsh4267):

A plane flies 420 miles with the wind and 340 miles against the wind in the same length of time. If the speed of the wind is 20 mph, what is the speed of the plane in still air?

OpenStudy (anonymous):

420/ (v +20) = 340/ ( v - 20) 420 * ( v - 20) = 340 (v +20) 80 v = 6800 + 8400 -> v = 15200 / 80 = 190 mph